(10%) Problem 2: Electrons have a charge of q--1.602 x10-19 C. A student observes that a Styrofoam cup has a total charge oi -0.205 uC 50% Part (a) write an expression for the number of excess electrons Nin the cup. Gr De HOME Att 1 2 3 deta 4 BACKSPACE CLEAR Submit Hint I give up! Hints: 0% deduction per hint. Hints remaining: 2 Feedback: 0% deduction per feedback. 50% Part (b) How many excess electrons are in the cup?
Answer #1

Given is:-

Charge on the electron  qー一1.602 × 10-19C

Total charge on the cup  Q=-0.205 × 10-6C

Now,

part-a

The expression for the total number of excess electrons N in the cup is

Total charge charge of each electron

or

oxed{N = rac{Q}{q}}

Part-b

Number of excess electron in the cup is

-0.205 × 10-6 -1.6 × 10-19 N=

which gives us

N = 1.28 × 1012| electrons
Answer #2
Go fuck yourself